OpenStudy (anonymous):

The manager of a mail order business will ship the packages to their customers. The cost, in dollars, to mail a package is based on its weight x, in pounds, and is given by the equation y = 0.7x2 –2.5x + 3. How much will the cost be to mail a 5 lb. package?

OpenStudy (anonymous):

Is 0.7x2 multiplying or is that a variable?

OpenStudy (anonymous):

I think its a variable

OpenStudy (anonymous):

Okay thanks, one moment

OpenStudy (anonymous):

y = 0.7x2 - 2.5 +3 y = 0.7 + 4.5 +3 y = 5.2 + 3 y = 8.2 I don't know if i'm right, if you go into the math chat room, you could ask for assistance.

OpenStudy (anonymous):

That's fine, Thanks for the help!

OpenStudy (anonymous):

Sure thing :P

OpenStudy (anonymous):

yeah it's 8.